### v1

# True Zero-Shot Inference of Dynamical Systems Preserving Long-Term Statistics

Abstract
Complex, temporally evolving phenomena, from climate to brain activity, are governed by dynamical systems (DS). DS reconstruction (DSR) seeks to infer generative surrogate models of these from observed data, reproducing their long-term behavior. Existing DSR approaches require purpose-training for any new system observed, lacking the zero-shot and in-context inference capabilities known from LLMs. Here we introduce DynaMix, a novel multivariate ALRNN-based mixture-of-experts architecture pre-trained for DSR, the first DSR model able to generalize zero-shot to out-of-domain DS. Just from a provided context signal, without any re-training, DynaMix faithfully forecasts the long-term evolution of novel DS where existing time series (TS) foundation models, like Chronos, fail – at a fraction of the number of parameters and orders of magnitude faster inference times. DynaMix outperforms TS foundation models in terms of long-term statistics, and often also short-term forecasts, even on real-world time series, like traffic or weather data, typically used for training and evaluating TS models, but not at all part of DynaMix’ training corpus. We illustrate some of the failure modes of TS models for DSR problems, and conclude that models built on DS principles may bear a huge potential also for advancing the TS prediction field.
1 Introduction
Figure 1: DynaMix achieves zero-shot DSR of attractor geometry and long-term temporal properties (power spectrum) from a short context signal while Chronos [2] fails.
Most real-world processes, from atmospheric phenomena and stock markets to brain activity or ecological networks, can be described as dynamical systems (DS) [54, 15, 82, 57]. Reconstructing these from observational data, called dynamical systems reconstruction (DSR), has been a long-standing challenge in scientific modeling [92, 23, 31]. DSR goes beyond conventional time series (TS) modeling, as we wish to have a generative model of the underlying process which exhibits long-term behavior with the same temporal and geometrical signatures (Fig. 1), i.e. the same invariant or ‘climate’ statistics [62, 66], as important in scientific applications. Achieving this usually requires special control-theoretic training techniques [56, 38] or loss objectives [65, 66, 40, 72] which accentuate the system’s long-term dynamics. Numerous deep learning approaches for DSR, based on recurrent neural networks (RNNs; [81, 22, 83, 16, 6, 69, 38, 7, 66]), neural ODEs [17, 43, 1, 45], Koopman operators [14, 53, 60, 3, 58, 28, 86], or library-based methods [12, 52, 41, 19, 55], have been advanced over the years. However, all of these require purpose-training on the specific system observed and struggle to generalize beyond their training distribution [33].
Inspired by the strong in-context and zero-shot generalization abilities of LLMs [10, 27, 21, 18], there has been a push recently to develop models with likewise properties for the time series domain. Time series foundation models like Chronos [2] or Mamba4Cast [5] are pre-trained on a huge database of real-world and artificial time series, and then tasked to forecast novel time series from which snippets are presented in-context, without any parameter fine-tuning. They are not built for DSR, however, and – as we show here – typically fail to properly capture a system’s long-term behavior and the structure of its attractors (Fig. 1; [93]).
To address this gap, we introduce DynaMix, a novel mixture-of-experts model designed for DSR and zero-shot forecasting. By training across a diverse range of DS, DynaMix learns transferable representations, enabling it to forecast previously unseen systems without retraining, including their underlying attractor geometries and invariant statistics (Fig. 1). The core features of DynaMix are:
  • Accurate zero-shot DSR: DynaMix achieves strong generalization across diverse DS, eliminating the need for extensive fine-tuning while maintaining accuracy in reproducing attractor geometry and long-term statistics.
  • Multivariate information transfer: Due to its multivariate architecture, the model efficiently captures dependencies among multiple system dimensions, enabling accurate reconstruction of their coupled dynamics. Further, it is neither bound to a specific dimensionality, but can flexibly adapt to other dimensions through specific embeddings, nor to a specific context length.
  • Computational and parameter efficiency: The model achieves high performance with a very lightweight architecture, enabling orders of magnitude faster inference than other foundation models.
  • Interpretable dynamics composition: The model provides insight into the dynamical composition of reconstructed systems, elucidating similarities between different DS.
  • Time series forecasting: Beyond DSR, our model is applicable to general time series forecasting where we employ different embedding techniques to obtain data representations accessible to DSR.
We evaluate our model on multiple benchmark DS and real-world time series, demonstrating superior zero-shot generalization on DSR problems compared to current TS foundation models.
2 Related work
Dynamical systems reconstruction (DSR)
In DSR we seek to learn generative models from time series data that represent the underlying system dynamics with all its topological, geometrical, and temporal properties [23, 33, 66, 30, 13, 11], in that sense providing an approximation to the underlying system’s governing equations. By definition, such a model should not only render viable short-term forecasts, but also reproduce the long-term evolution of the DS it has been trained on, both in state space and in the time domain. Methods approaching this goal have been founded on predefined function libraries, such as Sparse Identification of Nonlinear Dynamics (SINDy; [12, 52, 19, 41, 55]), on reservoir computers [63, 65, 66], neural ODEs [17, 43, 1, 45], Koopman operators [14, 53, 60, 3, 58, 28, 86], or different types of RNNs [81, 22, 83, 16, 6, 69, 38, 7]. More important than the architecture itself seems to be their proper training in order to ensure long-term (invariant) statistics of the underlying system are met: Methods like sparse [56, 6] or generalized [38] teacher forcing allow RNN-generated trajectories to ‘explore the future’ whilst training, yet keep loss gradients in check. Alternatively, long-term statistics based on the observed system’s Lyapunov spectrum, fractal geometry, or invariant measures may be added to regularize the loss [65, 66, 40, 72], but require to compute these properties from the data first. Despite these advances, out-of-domain generalization remains a key challenge in DSR [33]. Meta- and hierarchical learning models, trained across many DS simultaneously, have recently been designed as steps toward foundation models for DSR [9, 59, 44, 91], but still require parameter fine-tuning and lack in-context inference capabilities.
Time series foundation models
Large language models (LLMs) exhibit an impressive ability to infer patterns from prompts (context) and generalize to novel situations without retraining [10, 27, 21, 18], although the underlying reasons for this are still a matter of debate [49, 85, 90]. This inspired the development of general-purpose time series (TS) foundation models which could accurately forecast time series from a short segment provided ‘in-context’ [2, 26], without the need of task-specific fine-tuning [20]. One idea is to simply use pretrained LLMs directly as TS forecasters [34, 67, 75, 88, 79], but the inherent differences between textual and many temporal data, such as continuity, present significant challenges [78]. To overcome these, transformer-based architectures, similar in design to LLMs, such as Chronos [2] or Moirai [51], were specifically pretrained on a large corpus of time series data. Promising zero-shot forecasting capabilities have also been achieved with alternative architectural designs, such as Tiny Time Mixers [26] or state-space models like Mamba [5].
The success of TS foundation models raised hope these could also be utilized for zero-shot DSR, but the – to our knowledge – so far only previous study on this, based on Chronos, had mixed outcomes and fell short of a full DSR evaluation [93]: Successful DSR, if present at all, heavily depended on the initial conditions, and mechanisms such as context parroting visually gave the wrong illusion that features of the dynamics had been captured. In fact, here we show that existing TS foundation models are generally not capable of producing valid DSRs (Fig. 1), and highlight some of their failure modes. This is in contrast to the zero-shot DSR foundation architecture we develop here, based on models [7] and training algorithms [56] successful in DSR.
3 Methods
3.1 Model architecture
To enable zero-shot reconstruction of novel DS, here we develop a specific mixture-of-experts (MoE) architecture that can be pretrained across many diverse DS (Fig. 9), with different experts possibly specializing in different dynamical regimes. As a base model for the experts, we leverage a recent parameter-friendly RNN designed to yield topologically parsimonious and interpretable representations of DS it is being trained on, the Almost-Linear RNN (AL-RNN; [7]):
$\bm{z}_{t}=\bm{A}\bm{z}_{t-1}+\bm{W}\Phi^{*}(\bm{z}_{t-1})+\bm{h}\;.$ (1)
The model describes the evolution of an $M$-dimensional latent process $\bm{z}_{t}\in\mathbb{R}^{M}$, with linear self-connections $\bm{A}\in\text{diag}(\mathbb{R}^{M})$, weight matrix $\bm{W}\in\mathbb{R}^{M\times M}$, bias term $\bm{h}\in\mathbb{R}^{M}$, and $\Phi^{*}(\bm{z}_{t})$ defined as
$\Phi^{*}(\bm{z}_{t}):=\left[z_{1,t},\cdots,z_{M-P,t},\max(0,z_{M-P+1,t}),\cdots,\max(0,z_{M,t})\right]^{T},$ (2)
i.e. with a ReLU nonlinearity on only $P<<M$ out of the $M$ AL-RNN units. The first $N$ units are interpreted as the network’s readouts and provide the predicted observations, $\bm{\hat{X}}=\{\bm{\hat{x}}_{t}=\bm{z}_{1:N,t}\}\in\mathbb{R}^{N\times T}$, where $N$ is the observation dimension.
The selection of AL-RNN experts is achieved through a gating network (Fig. 2), which receives as inputs the (generally multivariate) context time series $\bm{C}=\{\bm{c}_{t}\}\in\mathbb{R}^{N\times T}$ as well as the current latent state $\bm{z}_{t}$. Both are passed into a state attention mechanism defined as
$\bm{w}^{att}_{t}=\sigma\left(\frac{\left|\bm{C}-\left(\bm{Dz}_{t}+\bm{\epsilon}\right)\bm{1}_{T}^{\top}\right|^{\top}\bm{1}_{N}}{T_{\text{att}}}\right)\in\mathbb{R}^{T}\;,$ (3)
where $\bm{D}\in\mathbb{R}^{N\times M}$ is a learnable matrix which maps the latent state into observation space, $\bm{\epsilon}\sim\mathcal{N}(0,\bm{\Sigma})$ exploration noise with a learnable covariance $\bm{\Sigma}$, $T_{\text{att}}$ a learnable temperature parameter, $\bm{1}_{\{T,N\}}$ column vectors of ones of length $T,N$, respectively, and $\sigma(\cdot)$ the softmax returning normalized weights. Hence, this mechanism computes attention weights based on some distance between projected latent states and actual context observations. At the same time, the context signal $\bm{C}$ is processed by a CNN, yielding temporal features $\tilde{\bm{C}}\in\mathbb{R}^{N\times T}$. These features are then weighted with the attention weights $\bm{w}^{att}_{t}$, and together with the current state $\bm{z}_{t}$ further processed by a multi-layer perceptron (MLP) followed by a softmax, yielding a set of expert weights at any time $t$:
$\bm{w}^{exp}_{t}=\sigma\left(\frac{\text{MLP}(\tilde{\bm{C}}{\bm{w}^{att}_{t}},\bm{z}_{t})}{T_{\text{exp}}}\right)\in\mathbb{R}^{J}\;,$ (4)
where $T_{\text{exp}}$ is another learnable temperature parameter. The forward-iterated latent states $\bm{z}_{t+1}^{j}$ of the individual experts $j\in\{1,...,J\}$ are then weighted by $\bm{w}^{exp}_{t}$ to yield the next time step prediction $\bm{z}_{t+1}=\sum_{j=1}^{J}w^{exp}_{j,t}\cdot\bm{z}_{t+1}^{j}$. Fig. 2 illustrates the whole approach. Note that one key advantage of this compared to other TS foundation architectures is that the context length is flexible by design, i.e. the CNN and the attention mechanism may take as input a context signal of arbitrary length.
Figure 2: Illustration of the DynaMix architecture. At test time, only a context signal $\bm{C}$ is provided and guides the selection of experts to yield arbitrarily long forward predictions of the dynamics.
3.2 Model training
For training, a collection of just $34$ different chaotic and cyclic DS is used [29], see Appx. A.2 for full specification. From each, multivariate time series $\bm{X}\in\mathbb{R}^{N\times T}$ are simulated ($\approx 1.6$ million in total) and then standardized (dimension-wise), of which the first $T_{C}<T$ column entries are defined as the context signal $\bm{C}=\bm{X}_{1:T_{C}}$. The experts are initialized at $t_{0}=T_{C}-\Delta t+1$ and forward-iterated until time $T$, such that they overlap for a period $\Delta t$ with the context. This is to ensure that the model learns to optimally utilize the context for generalization. A MSE loss is then computed across ground truth, $\bm{X}_{T_{C}-\Delta t+1:T}$, and respective model-generated time series $\hat{\bm{X}}$. The model is trained by sparse teacher forcing (STF; [6, 56]), a control-theoretic technique specifically designed for DSR. STF replaces part of the forward-iterated latent states $\bm{z}_{t}$ by data-inferred states $\hat{\bm{z}}_{t}$ at optimally chosen intervals $\tau$, to avoid exploding gradients even for chaotic systems while enabling the DSR model “to explore the future” (see [56] for detailed theoretical motivation). As shown in Fig. 24, STF with optimal $\tau$ is essential for achieving good zero-shot DSR results. Note that STF is only used for training and turned off at test time.
3.3 DSR evaluation & time series forecasting
We used two established measures to assess DSR quality [46, 6, 38, 89, 93, 66, 56, 61, 30]: First, to quantify agreement in state space (attractor) geometry, we employed a Kullback-Leibler divergence defined across space, $D_{stsp}$ (see Appx. A.1 for details). For assessing agreement in long-term temporal properties, we used the Hellinger distance $D_{H}$ defined across power spectra of the true and model-generated trajectories. In addition, a short-term, $n$-step-ahead mean absolute prediction error was evaluated. As has been emphasized repeatedly in the statistical and DSR literature [89, 46, 6, 56, 64, 30], prediction errors defined on time series are often only sensible on short time scales, because of the well known exponentially fast divergence of trajectories in chaotic DS (see Fig. 22).
Empirically observed time series are often just one-dimensional ($N^{*}=1$), but come from an inherently much higher-dimensional underlying DS. Our approach can in principle handle context signals of dimension $N^{*}<N$ smaller than those used in training by, e.g., just filling up $N-N^{*}$ dimensions in $\bm{C}$ with zeros. However, to work efficiently in a DSR setting, the context signal should represent the underlying DS as good as possible, for which it commonly needs to be lifted into a higher-dimensional space where sets explored by the dynamics ideally become diffeomorphic to those in the system’s true state space. Temporal delay embedding is the most popular technique to achieve this [76, 70]. For a $1d$ time series $\{x_{t}\}$, a $d$-dimensional embedding can be defined as [47]
$\bm{x}^{emb}_{t}=(x_{t},x_{t-\tau_{1}},...,x_{t-\tau_{d-1}})\;,$ (5)
where the $\tau_{i}$ are time lags, commonly estimated from the autocorrelation of the time series. Alternatively, if the delay embedding would become too large ($d>N$), we augment the empirically observed time series by a variant of the positional encoding common in transformers, defined here as
$\bm{x}^{emb}_{t}=\left(x_{t},\sin{\left(\frac{2\pi t}{\tau}+\phi_{1}\right)},...,\sin{\left(\frac{2\pi t}{\tau}+\phi_{N-1}\right)}\right)$ (6)
where $\tau:=\operatorname*{arg\,max}_{\tau>\tau_{min}}\mathbb{E}[x_{t}x_{t+\tau}]$ is given – provided it exceeds a threshold $\tau_{min}$ – by the maximal autocorrelation, and a random phase $\phi_{i}\in\left[0,\frac{\pi}{2}\right]$ is assigned to each dimension.
4 Results
4.1 Zero-shot DSR
Figure 3: a) DynaMix zero-shot DSR (red) compared to ground truth (blue) using a $500$-step context (gray) for the Sprott M system. b) Zero-shot forecasts (red) for the Selkov DS (true vector field in blue) from different initial conditions outside the context range (gray). c) DSR quality as a function of context length. d) DSR quality as a function of the temporal resolution $\Delta t$ of the context signal.
Figs. 1 (top) & 3a present example zero-shot DSRs of new chaotic DS, which were not part of DynaMix’ training domain, from just a provided context signal $\bm{C}$ (see Figs. 10-14 for further examples). Despite the inherent challenge of predicting chaotic systems, where small differences in initial conditions lead to rapidly diverging trajectories (Fig. 22), DynaMix successfully captures the system’s temporal & geometrical behavior, i.e. accurately reconstructs the underlying state space and faithfully recovers the true system’s power spectrum. Fig. 3b highlights further aspects of DynaMix’ reconstructions: Even though not trained on any $2d$ systems, the model successfully recreates the limit cycle from the $2d$ Selkov system [73]. Notably, the model even generalizes to new initial conditions outside the scope of the context data, i.e. correctly infers properties of the state space beyond the short context trajectory. As further shown in Fig. 3c, the context length necessary to achieve good DSR quickly converges, with as few as $500$ time steps often sufficient. Fig. 3d illustrates that our approach also works across a surprisingly wide range (about an order of magnitude) of temporal resolutions at which the underlying DS was sampled, confirming its robustness w.r.t. sampling frequency.
We next compared the zero-shot, out-of-domain DSR performance of DynaMix to that of a number of recent TS foundation models, including various versions of Chronos [2], Mamba4Cast [5], and Tiny Time Mixers [26], using the long-term statistics $D_{stsp}$ and $D_{H}$ for evaluation (see sect. A.1). As seen in Fig. 4a-b, DynaMix clearly and significantly outperforms all of these. In terms of short-term ($10$-step ahead) forecasts, for which TS foundation models (unlike DynaMix) are optimized, DynaMix performs on par with the (strongest in our batch) Chronos-t5 models (t-tests, all $p>0.37$), as shown in Fig. 4c. Table LABEL:tab:performance_DSR assembles further results on performance comparisons, while Figs. 15-20 provide specific examples. Moreover, a significant limitation of particularly the large Chronos models with millions of parameters is their extreme inefficiency in both inference time and computational costs, as compared in Fig. 5. In fact, their inference time can even exceed that of a custom-trained model. In contrast, DynaMix provides a highly efficient alternative, offering superior DSR performance while using much less parameters and computation time.
Figure 4: Zero-shot DSR performance across all 54 test set DS for DynaMix and various TS foundation models for context length $T_{C}=2000$ (see Fig. 21 for results with $T_{C}=512$). Median±$\pm$±MAD of a) $D_{stsp}$ (geometrical disagreement), b) $D_{H}$ (temporal disagreement), and c) short-term prediction error.
Figure 5: Inference times for zero-shot generation of Lorenz-63 DS with $10000$ forecasting steps. All models were provided the exact same context and run on the same hardware. Note the log-scale on the $x$-axis. For comparison, also the time needed for training and inference of a custom-trained AL-RNN is shown in red.
4.2 Reasons for the failure of TS foundation models on DSR
Why do TS foundation models perform so poorly on DSR problems? Essentially, they face three key issues: First, most TS foundation models cannot efficiently deal with multivariate time series, but treat dimensions independently. However, in a nonlinear DS all variables are usually coupled, and their joint evolution is actually defining for the system’s (long-term) dynamics and attractor states. Thus, ignoring this, as illustrated in Fig. 6a, results in poor long-term accuracy (see also Figs. 15-20). This highlights a fundamental limitation: Univariate approaches are inherently insufficient for reconstructing multivariate, interacting DS. Second, TS foundation models are trained for short-term prediction, but not for DSR, providing another reason why they fail to reconstruct the correct long-term behavior. As illustrated in Figs. 1 & 6b, in the longer run they often converge to fixed points or simple cycles where the true dynamics is chaotic, not reflecting the context anymore. Third, context parroting, a phenomenon originally identified in LLMs [4], has also been observed in TS foundation models [93], where they tend to repetitively reproduce the exact input pattern rather than dynamically adapting to the system’s true evolution. Thus, TS foundation models simply produce oscillations, where the true dynamics is chaotic and therefore by definition irregular and non-repeating (Fig. 6c). This is illustrated in more detail in Fig. 6d for Chronos-t5-base, the best of the TS foundation models in our batch. It is particularly evident in the power spectrum (Fig. 6d right), where Chronos’ sharp, narrow peaks wrongly suggest periodicity, while the true spectrum is usually rather broad and smeared out for chaotic systems, as the one exhibited by DynaMix. We further quantified this by calculating from the forecast time series the maximum Lyapunov exponent [56], a measure for the exponential divergence rate of trajectories (see Appx. A.1). For Chronos, the value $\lambda_{Chronos}\approx 0.06$ close to zero confirms the nearly periodic behavior, while DynaMix produces exponentially diverging trajectories with $\lambda_{DynaMix}\approx 1.48$ close to the ground truth value of $\lambda_{Lorenz63}\approx 1.37$. Hence, unlike DynaMix, Chronos does not reconstruct the underlying dynamics but just repeats the context. On the other hand, if the underlying dynamics truly is cyclic, DynaMix settles into a cyclic pattern as well, exhibiting no intrinsic bias, as illustrated in Fig. 6e.
Figure 6: Common problems in TS foundation models (red: forecast, gray: context, blue: ground truth). a) Dimensions are dynamically decoupled. b) Long-term forecasts often converge to simple fixed points or cycles unrelated to the true dynamics. c) Context parroting covers up the true dynamics. d) Example forecasts for the chaotic Lorenz-63 system with DynaMix and Chronos-t5-base. Chronos’ context parroting results in cyclic repetition of a fixed pattern, thus inherently missing the true aperiodic, chaotic dynamics. This is evident in the close-to-0 max. Lyapunov exponent (center) and peaked power spectrum (right). e) For comparison, if the underlying behavior is truly cyclic with prominent peaks in the power spectrum, both Chronos and DynaMix are able to capture it.
4.3 Dynamical similarity
DynaMix’ mixture-of-experts design produces a natural similarity measure for comparing the dynamics of different systems, by evaluating the relative contribution of each of the experts to the forecast dynamics. As shown in Fig. 7a (bottom), different experts specialize in specific dynamical patterns or regimes. Hence, we can utilize the average expert usage $\bm{e}=\frac{1}{T}\sum_{t=1}^{T}\bm{w}^{exp}_{t}$ across the generated time series to construct a similarity metric
$S=\frac{1}{S_{max}}\left(||\bm{e}_{1}-\bm{e}_{2}||_{1}\right)^{-1}\;,$ (7)
defined here as the inverse $L_{1}$ norm of the $\bm{e}$ vectors between any two systems, rescaled to the range $[0,1]$.
Figure 7: a) Top: Vector fields of different experts which in combination determine the attractor. Bottom plot illustrates the dominant expert in each time window. b) Similarity matrix among all $84$ types of attractor dynamics in our dataset, determined by computing $S$ from the reconstructed time series. c) Examples of two pairs of attractors classified as similar (top), and one dissimilar pair.
4.4 Time series forecasting from a DS perspective
Figure 8: Time series forecasts for different data domains (gray: context, blue: ground truth, other colors: various model predictions): Forecasts of a) partially ($1$d) observed Lorenz-63 DS, b) electricity transformer temperature data, c) hourly car traffic data with weekly cycle, d) Huawei cloud request data, e) soil temperature development, f) air pressure data, g) human functional magnetic resonance imaging (fMRI) data, h) human electroencephalogram (EEG) data.
Table 1: Performance comparison on empirical time series in terms of geometrical divergence ($D_{\text{stsp}}$), long-term temporal distance ($D_{H}$), and forecast error (MAE). Best in red, second-best in blue.
System DynaMix Chronos-t5-base Mamba4Cast TTM
$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
Partially obs. DS 0.11 0.25 0.32 0.11 0.32 0.39 8.14 0.35 0.76 6.54 0.46 0.89
ETTh1 0.8 0.08 5.97 0.33 0.09 3.22 0.94 0.1 3.63 6.15 0.09 4.64
Traffic 0.68 0.21 8.42 0.35 0.1 3.16 1.4 0.34 7.68 1.77 0.25 5.25
Cloud Requests 0.17 0.14 147.28 1.57 0.3 557.35 7.91 0.16 282.95 1.39 0.26 147.54
Weather (Temp.) 0.59 0.09 2.24 4.72 0.22 4.14 6.96 0.58 9.39 5.39 0.3 2.76
Weather (Press.) 0.41 0.19 8.85 6.62 0.23 9.05 7.03 0.52 8.6 7.27 0.25 10.03
Human fMRI 0.63 0.09 0.56 5.77 0.12 0.85 7.99 0.11 1.35 6.66 0.21 0.9
Human EEG 0.78 0.23 1.09 10.17 0.24 1.1 6.0 0.62 1.12 2.48 0.37 0.92
We next tested and compared DynaMix on a variety of real-world datasets often used to probe time series models. As discussed in sect. 3.3, for DynaMix we used a multivariate embedding of the typically $1d$ empirical time series, to better reflect the underlying DS properties. Fig. 8a confirms for a $1d$ time series from the $3d$ chaotic Lorenz-63 system that this works well and DynaMix, in contrast to all other models, is still able to reconstruct the chaotic behavior. When applied to real-world time series, such as traffic, weather, or cloud request data, where previous TS foundation models have shown inconsistent results [80], our model produces accurate forecasts which reflect the dynamics well (Fig. 8b-f, see Appx. A.2 for further details on the datasets). Our model seems to particularly excel on human physiological signals, such as functional magnetic resonance imaging (fMRI) or electroencephalographic (EEG) recordings, where all other models essentially produce meaningless dynamics, at best capturing the mean trend (Fig. 8g,h).
Table 1 quantitatively confirms that DynaMix mostly outperforms the other models tested on the DSR measures. Surprisingly, it often even outperforms them in terms of the forecasting error MAE (mean-absolute error), although its training corpus consisted purely of the simulated DS shown in Fig. 9, i.e. did not include any empirical data at all, let alone traffic, cloud usage, electricity, or temperature data! This is in stark contrast to the other TS foundation models tested here, which mostly had examples of such empirical data in their training corpus and hence might be expected to have an edge over DynaMix. Many real-world data, like weather, human fMRI, or EEG data, are known to bear signatures of deterministic chaos [56, 48], potentially at least partly explaining DynaMix’ strongly competitive forecasts, although it has never seen such type of data in training.
5 Conclusions
Here we introduce the first DSR foundation architecture that achieves zero-shot reconstructions of novel DS, accurately reproducing their long-term statistics, just from a provided context signal without any retraining or fine-tuning. It hugely outperforms current TS foundation models like Chronos in terms of DSR quality, although much more lightweight with only a fraction (0.1%) of the number of parameters. Consequently, DynaMix features inference times about an order of magnitude faster than the closest competitor. Most surprisingly, it even outperforms current TS foundation models on empirical data of types it has – in contrast to its competitors – never seen in training, often producing also better short-term forecasts. All of this is achieved with a fairly narrow training corpus consisting mainly of $3d$ synthetic data from chaotic systems, which suggests there may be considerable room for further improvement by extending DynaMix’ training to include, e.g., various empirical time series.
One important take-home from this work therefore is that foundation models built based on principles of DS theory may be able to profoundly improve performance of current zero-shot TS forecasters. Most (if not all) empirically observed time series come from some underlying DS, and acknowledging this fact in model training and construction may help to advance the field. Besides DynaMix’ specific DSR architecture, control-theoretically motivated training techniques, and embedding principles, also the fact that it has seen many chaotic systems in training may play a role. Most complex DS in nature [15, 82, 57], engineering [74], and society [54, 87] are likely chaotic, featuring an infinite spectrum of potential temporal patterns (unstable periodic orbits; [36]).
Limitations
For now, we have mainly focused on rather stationary synthetic and empirical time series (but see Fig. 8b). Changes in statistical properties over time, discontinuities, or widely differing time scales in the data observed, impose severe difficulties for zero-shot forecasting, as illustrated in Fig. 23a,b, which also plague other TS foundation models. However, this may partly be amended by explicitly including non-stationary DS or DS evolving on many different time scales, featuring sharp transitions, etc., in the training corpus. Incorrect embeddings, which may not be immediately obvious, can also lead to forecast failures (Fig. 23c).
Acknowledgements
This work was funded by the German Research Foundation (DFG) within Germany’s Excellence Strategy EXC 2181/1 – 390900948 (STRUCTURES), and through DFG individual grant Du 354/15-1 to DD.
Appendix A Appendix
A.1 Methodological details
Training method
For training our model, we used a variant of sparse teacher forcing (STF), a control-theoretic method for DSR, with the theoretical background and rationale developed in [56]. STF and related training methods [6, 8, 38, 56, 24] circumvent the exploding gradient problem when training on chaotic systems, while still enabling the DSR model to sufficiently forward-iterate trajectories into the future to capture long-term properties (see [56] for details), and produces state-of-the-art results for DSR [7, 9]. In STF, current latent states states $\bm{z}_{t}$ are replaced with states inferred from the data at fixed intervals $\tau$ by (pseudo-)inversion of the decoder model, $\tilde{\bm{z}}_{t}=g^{-1}(\bm{x}_{t})$, thus recalibrating the trajectory at times ideally chosen based on the system’s Lyapunov spectrum [56] (or simply treating $\tau$ as a hyper-parameter). In our case of an identity mapping from a latent subspace to the observations, $\hat{\bm{x}}_{t}=\mathcal{I}_{(N\times M)}\bm{z}_{t}$, this inversion becomes trivial, where $N$ is the number of observations (and thus readout neurons) and $\mathcal{I}_{(N\times M)}$ an $N\times M$ matrix with $\mathcal{I}_{(N\times M),rr}=1$ for $r\leq N$ and zeros else. Thus, during training, we force
$\bm{z}_{t+1}=\begin{cases}F_{\bm{\theta}}(\tilde{\bm{z}}_{t},\bm{C})&\text{if }t\in\mathcal{T}=\{l\tau+1\}_{l\in\mathbb{N}_{0}}\\ F_{\bm{\theta}}(\bm{z}_{t},\bm{C})&\text{otherwise, }\end{cases}$ (8)
where $\tilde{\bm{z}}_{t}=(\bm{x}_{t},\bm{z}_{N+1:M,t})^{T}$, $F_{\theta}$ is the mixture of experts using the context $\bm{C}$, and $\tau=10$ here (see Fig. 24). Importantly, STF is only used for training the model, applied after calculating the loss, and is turned off at test time.
For the loss, we simply use the standard mean squared error (MSE) between model predictions $\hat{\bm{X}}$ and ground truth observations $\bm{X}$,
$\mathcal{L}_{MSE}(\hat{\bm{X}},\bm{X})=\frac{1}{N\cdot(T-T_{C}+\Delta t)}\sum_{t=T_{C}-\Delta t+1}^{T}\left\lVert\hat{\bm{x}}_{t}-\bm{x}_{t}\right\rVert_{2}^{2}.$ (9)
To this we add a regularization term meant to encourage the model to explore a wider region of state space based on the context, by enhancing the variance $\bm{\Sigma}$ of the exploration noise in the state attention:
$\mathcal{L}_{reg}=\lambda\left(1+\sum_{i}\Sigma_{ii}\right)^{-1}\;,$ (10)
where we chose $\lambda=0.01$. Rectified adaptive moment estimation (RADAM) [50] was employed as the optimizer, with $L=50$ batches of $S_{B}=16$ sequences per epoch, each of length $T=550$, and $2000$ epochs in total. We used a learning rate exponentially decaying from $\eta_{\text{start}}=2\cdot 10^{-3}$ to $\eta_{\text{end}}=10^{-5}$. The context length used in training was set to $T_{C}=500$, and the window of overlap with the model-generated time series to $\Delta t=50$, see sect. 3.2. Training was performed on a single CPU (18-Core Xeon Gold 6254), and a single epoch took $60$ seconds depending on sequence length and model size. At test time, no retraining or fine-tuning is performed, but utilizing the context $\bm{C}$ the model just forward-iterates from the last context time step. Hyperparameters for training were partly selected according to previous results in [7, 6], and partly a few different parameter settings were tested, namely $\lambda=\{0.0,0.01,0.02\}$, $\Delta t=\{0,50,100\}$, and $T_{C}=\{250,500\}$, as extensive grid search was not necessary to obtain a well-performing model.
Model details
We use $J=20$ AL-RNN experts for our model. Each expert has a latent dimension of $M=20$, of which $P=7$ are rectified-linear units (ReLUs). As suggested in [7], we only use the first $N$ of the $M-P$ linear units for the readout. We followed the initialization protocol in [6, 77], drawing $\bm{W}$ from a Gaussian with mean $\bm{0}$ and $\bm{\sigma}=0.01\mathds{1}$, setting $\bm{h}=\mathbf{0}$, and $\bm{A}$ to the diagonal of a normalized positive-definite random matrix. The initial latent state was estimated as
$\bm{z}_{1}=\begin{bmatrix}\bm{x}_{1}\\ \bm{Lx}_{1}\end{bmatrix},$ (11)
where $\bm{L}\in\mathbb{R}^{(M-N)\times N}$ is jointly learned with other model parameters.
[…]
Performance measures
As in [46, 56, 38, 8, 61, 30, 93], we assess the geometric (dis)similarity between true and model-generated attractors using a Kullback-Leibler divergence ($D_{\text{stsp}}$) defined across state space. Specifically, $D_{\text{stsp}}$ determines the overlap between the distributions of true trajectory points, $p_{true}(\bm{x})$, and of model-generated trajectories, $p_{gen}(\bm{x}|\bm{z})$, given by
$D_{\text{stsp}}(p_{true}(\bm{x})\lVert p_{gen}(\bm{x}|\bm{z}))=\int p_{true}(\bm{x})\log\frac{p_{true}(\bm{x})}{p_{gen}(\bm{x}|\bm{z})}d\bm{x}.$ (12)
While especially for higher-dimensional state spaces these distributions may be approximated by Gaussian mixture models [46], here a discretized version was sufficient, with the state space parcellated into $K=m^{N}$ bins, where $m$ is the number of bins per dimension and $N$ the dimensionality of the system. We estimate the occupation probabilities $p_{i}$ of each bin via the relative frequencies $\hat{p}_{i}$ of trajectory visits and approximate $D_{\text{stsp}}$ as
$D_{\text{stsp}}\approx\sum_{i=1}^{K}\hat{p}_{true;i}\log\frac{\hat{p}_{true;i}}{\hat{p}_{gen;i}}.$ (13)
We set $m=30$ bins per dimension, following [37]. To ensure a steady-state distribution is reached, long trajectories of $T=10,000$ time steps are sampled.
To evaluate the agreement in long-term temporal dynamics, we compute the Hellinger distance $D_{H}$ between the power spectra of the true and model-generated time series [56, 38], defined as
[…]
where $F(\omega)$ and $G(\omega)$ are the power spectra of the true and generated time series, respectively. Power spectra are obtained via the dimension-wise Fast Fourier Transform (FFT), smoothed using a Gaussian kernel (with $\sigma=20$ for DS and $\sigma=2$ for the short empirical time series), and normalized to ensure comparability. High-frequency tails dominated by noise are truncated as described in [38]. The aggregated Hellinger distance $D_{H}$ is then computed as the average across all dimension-wise spectral comparisons.
The $n$-step ahead prediction error to assess short-term forecast quality (with $n=10$ in our evaluation) is given by
$\text{PE}(n)=\left\|\bm{x}_{T_{C}+n}-\hat{\bm{x}}_{T_{C}+n}\right\|_{1}.$ (15)
The MAE for evaluating short term predictions on the empirical time series [39] is given by
$\text{MAE}=\frac{1}{n}\sum_{t=T_{C}+1}^{T_{C}+n}\left|x_{t}-\hat{x}_{t}\right|\;,$ (16)
where we chose $n$ according to the data’s temporal scale and resolution ({ETTh1,fMRI}: $n=40$,{traffic, cloud requests, partially obs. DS}: $n=80$, {human EEG, weather air pressure}: $n=120$, {weather temperature}: $n=200$).
Lyapunov exponent
For numerically estimating the maximum Lyapunov exponent $\lambda_{\text{max}}$ from a one-dimensional empirical time series $\{x_{t}\},t=1\dots T$, one common approach is the Rosenstein algorithm [68]. First, the time series is embedded into a state space using time-delay embedding as in eq. 5 (here using the PECUZAL algorithm [47]). For each embedded point $\bm{x}_{i}$, the nearest neighbor $\bm{x}_{j}$ is found under the constraints $|i-j|>l_{t}$, with $l_{t}$ a threshold to remove purely temporal neighbors (living on the same piece of trajectory), and a minimal initial state space separation $\|\bm{x}_{i}-\bm{x}_{j}\|_{2}>l_{s}$, where we chose $l_{t}=150$ (according to the mean periodicity [68]) and $l_{s}=0.2$. The local divergence between initially nearby trajectories is then tracked over a range of time steps $k=0\dots k_{max}$,
$d_{i}(k)=\|\bm{x}_{i+k}-\bm{x}_{j(i)+k}\|_{2}\;.$ (17)
The maximum Lyapunov exponent is given by
$d_{i}(k)\approx d_{i}(0)e^{\lambda_{\text{max}}k\Delta t}\;,$ (18)
where $\Delta t$ is the temporal resolution of the series. Taking the logarithm and averaging over all pairs leads to
$\langle\ln d(k)\rangle\approx\lambda_{\max}k\Delta t+C.$ (19)
In chaotic systems, for a suitable choice of $k_{max}$ (here: $50$), this quantity initially grows approximately linearly with time [42], with the slope given by $\lambda_{\max}>0$. In contrast, periodic (cyclic) systems have $\lambda_{\max}=0$.
A.2 Datasets
Training data
DynaMix is trained on about $1.6$ million simulated time series of length $T=550$ sampled from $34$ different $3d$ DS with cyclic or chaotic attractors, collected in [29]. Gaussian noise of $5\%$ of the data standard deviation was added to all dimensions, and time series were standardized to ensure comparable scaling of all data. Attractor dynamics were chosen to reflect different types of behavior, see Fig. 9.
Figure 9: Training data (blue) and their reconstruction in red.
Test data
Our test set for DS consists of simulated time series of length $10^{5}$ sampled from $54$ different $3d$ DS collected in [29], which are not part of the training set. Furthermore, we evaluated the Selkov 2d DS [73] describing a kinetic model of an open monosubstrate enzyme reaction given by the equations
$\frac{dx}{dt}$ $=-x+ay+x^{2}y,$ (20)
$\frac{dy}{dt}$ $=b-ay-x^{2}y,$
where we chose $a=0.1$ and $b=0.5$.
For evaluation of time series forecasting, different real-world data were used. The traffic data are hourly recordings of the number of cars passing road junctions (https://www.kaggle.com/datasets/fedesoriano/traffic-prediction-dataset/data). The cloud data, also used in [80] to evaluate TS foundation models, are publicly available from Huawei Cloud https://github.com/sir-lab/data-release. It consists of function requests from Huawei’s serverless cloud platform. The weather data consists of daily sampled soil temperature and air pressure measured in the city of Mannheim, Germany (source: Deutscher Wetterdienst [German weather service]), and can be accessed via https://www.dwd.de/EN/ourservices/cdc/cdc_ueberblick-klimadaten_en.html. The functional magnetic resonance imaging (fMRI) data comes from human subjects performing cognitive tasks and is publicly available on GitHub [48]. We followed Kramer et al. [48] and selected the first principal component of BOLD activity in each of the $20$ brain regions. Kramer et al. [48] report a positive maximum Lyapunov exponent for models reconstructed from these time series, indicating their chaotic nature (see also [84]). The ETTh1 dataset is part of the Electricity Transformer Temperature (ETT) benchmark, which is widely used for evaluating TS forecasting models. It contains hourly data collected from a power transformer station [94] and can be accessed at https://github.com/zhouhaoyi/ETDataset. Electroencephalogram (EEG) data were taken from a study by Schalk et al. [71], comprising 64-channel data collected from human subjects performing various motor and imagery tasks. Following the approach of Brenner et al. [6], the signals were smoothed using a Hann window of length $15$. As for the training data, all test data were standardized for processing by the models.
A.3 TS foundation models
Chronos
Chronos is a recent TS foundation framework that adapts a transformer-based LLM architecture for probabilistic time series forecasting [2]. Key to this approach is the tokenization of real-valued time series observations through scaling and quantization, transforming them into token sequences accessible to language models. The model is pretrained on extensive datasets, including both synthetic data generated via Gaussian processes, as well as an extensive batch of empirical time series from [2, 32], including traffic, weather/ climate, electricity and web data, enabling it to achieve strong zero-shot performance across diverse datasets without task-specific fine-tuning. For our evaluation we used the standard pipeline as described in https://github.com/amazon-science/chronos-forecasting.
Mamba4Cast
Mamba4Cast is a zero-shot time series forecasting model based on the Mamba architecture [35], a type of linear (‘state-space’) RNN with nonlinear input & output gating, and inspired by Prior-data Fitted Networks (PFNs) [5]. Trained exclusively on synthetic data, Mamba4Cast can generate zero-shot forecasts when provided with time series context information. For evaluation we follow https://github.com/automl/Mamba4Cast.
Tiny Time Mixers
Tiny Time Mixers is another time series foundation model using a non-transformer architecture [26]. It builds upon the TSMixer backbone [25] to which it adds several key innovations such as adaptive patching, diverse resolution sampling, and multi-resolution prefix tuning to improve generalization across datasets with varying temporal resolutions. Evaluation here is performed as in https://github.com/glehet/TTM1.
For comparability, all models were evaluated on the exact same CPU (18-Core Xeon Gold 6254) and GPU (Nvidia RTX 2080 Ti) using 512GB of RAM.
A.4 Further results
Figure 10: Zero-shot DSR (red) from a $2000$-step context of unseen DS not contained in the training corpus (ground truth in blue).
Figure 11: Zero-shot DSR of chaotic Lorenz-63 system (gray: context, blue: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
Figure 12: Zero-shot DSR of Lorenz-63 system in cyclic regime (gray: context, blue: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
Figure 13: Zero-shot DSR of chaotic finance system (gray: context, blue: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
Figure 14: Zero-shot DSR of chaotic Genesio Tesi system (gray: context, blue: ground truth, red: model-generated). Left: State space, center: time graphs, right: power spectrum.
Figure 15: Comparison of zero-shot forecasting of finance system for DynaMix vs. different TS foundation models.
Figure 16: Comparison of zero-shot forecasting of Sprott D system for DynaMix vs. different TS foundation models.
Figure 17: Comparison of zero-shot forecasting of Sprott M system for DynaMix vs. different TS foundation models.
Figure 18: Comparison of zero-shot forecasting of Vallise El Nino system for DynaMix vs. different TS foundation models.
Figure 19: Comparison of zero-shot forecasting of Sprott C system for DynaMix vs. different TS foundation models.
Figure 20: Comparison of zero-shot forecasting of Sprott A system for DynaMix vs. different TS foundation models.
Figure 21: Zero-shot DSR performance across all 54 test set DS for DynaMix and various TS foundation models as in Fig. 4, but for context length $T_{C}=512$. Detailed values in Table LABEL:tab:performance_DSR_CL512.
Figure 22: Exponential divergence of two initially nearby trajectories (in blue and red) illustrated for the chaotic Lorenz-63 system. The prediction error is still sensible on a short time scale ($\text{MAE}=0.09$), but then rapidly increases and breaks down as a suitable metric in the longer-term ($\text{MAE}=1.12$), although both trajectories were drawn from the exact same system with the very same parameters.
Figure 23: Failure of proper zero-shot DSR. a) DynaMix struggles with sharp peaks in an otherwise much more slowly evolving time series. b) As quantified in Fig. 3, selecting a too broad temporal resolution of the context time series can lead to improper reconstructions with, in this case, temporal disalignment w.r.t. the true signal. However, this is a general problem for any TS foundation model, and Chronos fares even much worse in this example. c) Zero-shot DSR may fail (red) as well if the context time series is not properly embedded. In this example, the required embedding dimension exceeded the model capacity. However, as shown in orange, this can be amended by using the positional encoding, eq. 6.
Figure 24: DSR performance for different teacher forcing intervals $\tau$, illustrating an optimal value is essential for successful training, see [56, 6].
Table 2: Performance comparison for zero-shot inference across all DS in the dataset from [29] using $T_{C}=2000$. Best value in red, second best in blue.
System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-t5-tiny Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-bolt-tiny Mamba4Cast TTM
$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
Atmospheric Regime 9.7 0.6 0.3 9.2 0.4 0.2 10.5 0.4 0.2 12.0 0.5 0.1 13.3 0.5 0.0 13.7 0.5 0.3 11.6 0.9 0.3 12.3 0.8 0.2 12.3 0.9 0.3 11.7 0.7 0.3 11.2 0.8 0.6 13.9 0.4 1.3
BeerRNN 8.2 0.4 3.6 12.1 0.5 0.8 9.3 0.4 1.3 13.3 0.4 1.6 10.9 0.3 1.3 13.1 0.5 1.5 12.4 0.9 1.7 12.7 0.8 2.1 11.9 0.9 2.0 11.7 0.8 1.9 11.4 0.8 1.1 14.2 0.4 2.5
Bouali 4.7 0.1 0.7 7.0 0.3 0.5 12.1 0.2 0.6 11.6 0.1 0.6 13.1 0.2 0.3 13.5 0.4 1.5 11.2 0.7 2.2 12.0 0.6 3.4 12.3 0.5 2.2 9.8 0.5 1.5 12.9 0.7 2.6 14.0 0.3 2.5
Burke Shaw 3.4 0.1 0.2 8.0 0.1 0.3 5.6 0.1 0.2 5.1 0.2 0.1 11.1 0.2 0.4 13.2 0.6 1.0 11.0 0.6 0.9 11.0 0.7 0.5 11.7 0.8 0.2 10.0 0.7 1.1 10.2 0.7 0.6 12.9 0.4 0.8
Ca Two Plus Quasip. 3.3 0.1 0.3 9.6 0.5 1.2 9.5 0.3 0.3 12.6 0.2 0.9 12.7 0.4 0.6 13.2 0.4 1.6 9.4 0.7 0.8 9.4 0.6 1.5 9.6 0.7 0.8 11.5 0.5 0.9 11.1 0.6 2.0 13.8 0.2 5.2
Chen Lee 5.1 0.4 0.1 4.2 0.2 0.0 3.6 0.2 0.0 3.9 0.2 0.0 5.3 0.3 0.0 12.0 0.8 0.3 4.5 0.8 0.3 8.8 0.7 1.1 7.7 0.7 0.5 7.2 0.7 0.5 11.3 0.8 0.2 8.9 0.4 0.6
Chua 3.0 0.1 0.4 12.9 0.2 0.4 12.4 0.4 0.4 13.3 0.2 0.1 13.5 0.4 0.4 13.5 0.3 0.9 11.9 0.6 1.0 11.7 0.4 0.6 11.7 0.4 0.6 12.0 0.4 0.7 12.1 0.4 0.2 13.4 0.3 1.0
Coullet 0.8 0.0 0.3 0.6 0.0 0.1 6.3 0.0 0.1 3.9 0.1 0.1 13.7 0.3 0.1 14.6 0.3 0.6 11.4 0.7 0.6 12.8 0.7 0.4 11.8 0.6 0.7 10.7 0.6 1.1 13.3 0.6 0.7 15.0 0.2 1.8
Dadras 3.0 0.2 13.9 10.0 0.5 9.2 9.2 0.5 4.2 11.8 0.4 7.5 11.4 0.3 4.7 12.9 0.8 9.0 10.5 0.8 8.9 9.4 0.7 9.0 11.0 0.6 8.8 8.8 0.5 8.9 13.1 0.8 7.9 12.2 0.4 9.2
Dequan Li 10.0 0.5 0.1 12.6 0.5 0.1 13.2 0.4 0.1 13.1 0.3 0.1 12.9 0.4 0.5 13.1 0.3 0.4 12.4 0.4 0.4 12.7 0.4 0.3 12.2 0.4 0.3 12.6 0.4 0.6 12.3 0.4 1.1 13.0 0.4 0.4
Finance 3.0 0.2 0.6 8.5 0.4 0.2 9.4 0.6 0.3 13.4 0.3 0.5 12.2 0.4 0.3 13.6 0.5 0.3 12.0 0.7 1.3 11.7 0.6 2.3 12.4 0.7 1.3 11.8 0.6 1.4 12.6 0.8 2.3 14.0 0.3 2.1
Genesio Tesi 10.5 0.2 0.3 8.9 0.1 0.1 8.7 0.1 0.0 7.7 0.1 0.0 11.8 0.4 0.2 12.3 0.3 0.3 10.7 0.3 0.2 11.4 0.9 0.8 10.6 0.6 1.1 10.0 0.4 1.6 10.4 0.8 0.4 12.3 0.2 0.7
Guckenheimer Holmes 7.9 0.0 0.2 10.6 0.4 0.1 11.7 0.2 0.0 11.3 0.4 0.1 13.2 0.7 0.1 12.5 0.2 0.1 10.7 0.6 0.1 10.6 0.6 0.2 9.2 0.6 0.1 10.3 0.6 0.1 10.4 0.5 0.5 11.6 0.3 0.6
Hadley 8.4 0.3 0.2 10.8 0.5 0.2 9.3 0.6 0.2 11.9 0.4 0.2 12.5 0.4 0.2 12.5 0.3 0.4 10.9 0.7 0.4 10.7 0.7 1.0 9.5 0.6 0.3 10.2 0.4 0.4 11.7 0.7 0.8 12.5 0.3 1.6
Halvorsen 2.8 0.1 0.2 6.0 0.1 0.2 8.4 0.1 0.2 5.9 0.1 0.3 12.6 0.2 0.3 13.3 0.3 1.3 11.8 0.8 1.1 9.5 0.7 0.8 8.9 0.8 0.4 10.4 0.5 0.4 12.2 0.6 1.3 13.2 0.3 1.5
Hindmarsh Rose 6.8 0.5 0.6 9.7 0.3 0.1 6.7 0.5 0.3 8.4 0.1 0.1 11.5 0.4 0.7 13.8 0.6 1.0 12.7 0.7 1.3 12.5 0.7 1.5 13.9 0.6 1.1 13.4 0.7 1.1 13.9 0.6 0.7 14.0 0.2 1.6
Isothermal Chemical 2.5 0.2 0.1 3.7 0.1 0.0 7.9 0.3 0.0 9.7 0.3 0.0 9.0 0.3 0.0 13.8 0.5 0.2 7.9 0.8 0.3 5.9 0.8 0.2 8.4 0.8 0.2 8.5 0.8 0.2 12.8 0.8 0.5 15.6 0.2 2.8
Itik Banks tumor 9.1 0.2 1.2 10.0 0.6 0.5 8.4 0.5 0.3 10.5 0.4 0.3 12.8 0.3 0.2 12.1 0.7 0.4 9.2 0.7 1.2 9.5 0.7 1.2 9.7 0.8 1.2 9.6 0.7 1.0 13.3 0.9 0.5 13.2 0.3 1.7
Jerk Circuit 2.7 0.3 0.4 10.1 0.4 1.0 9.7 0.4 0.4 7.5 0.3 0.4 11.1 0.4 0.4 12.3 0.3 0.2 10.8 0.8 1.6 10.9 0.5 2.0 10.5 0.9 1.9 10.1 0.9 1.8 11.1 0.9 1.8 13.9 0.4 2.1
Laser 6.5 0.3 0.1 11.1 0.4 0.1 10.8 0.2 0.1 12.8 0.2 0.2 13.8 0.5 0.4 14.1 0.5 0.3 11.8 0.7 0.3 11.5 0.7 0.4 11.0 0.7 0.5 11.9 0.7 1.6 13.3 0.6 0.7 13.8 0.2 1.2
Lorenz 1.1 0.1 0.7 6.6 0.2 0.3 6.0 0.2 0.3 10.0 0.2 0.2 12.0 0.4 0.3 13.0 0.5 0.3 10.9 0.7 0.3 11.3 0.5 0.6 12.7 0.6 0.9 12.8 0.3 1.8 11.4 0.7 0.2 13.0 0.3 1.9
Lorenz 84 6.9 0.3 4.9 9.5 0.4 5.2 11.0 0.4 4.8 11.5 0.4 3.2 12.2 0.4 2.5 12.9 0.5 4.4 12.9 0.8 2.4 12.9 0.8 2.5 11.5 0.9 2.4 12.3 0.7 2.4 13.4 0.8 3.3 13.4 0.5 2.4
Lu Chen 3.4 0.5 0.8 6.3 0.4 0.2 5.3 0.3 0.3 8.0 0.1 0.2 7.1 0.1 0.2 4.9 0.2 0.1 11.6 0.9 1.3 11.9 1.0 1.2 11.3 1.0 1.9 11.4 1.0 1.4 9.7 0.8 0.7 13.7 0.4 3.4
Moore Spiegel 5.9 0.4 1.0 10.4 0.6 1.0 10.0 0.5 0.0 12.5 0.4 0.2 9.9 0.5 0.6 12.1 0.3 0.9 10.5 0.9 2.8 11.4 0.9 1.8 10.1 0.9 2.5 9.8 0.7 3.4 9.4 0.8 4.6 12.5 0.3 3.6
Multi Chua 7.8 0.3 0.8 9.3 0.5 0.0 9.3 0.2 0.2 7.1 0.4 0.1 13.3 0.2 0.4 14.0 0.3 1.2 11.5 0.6 1.8 12.2 0.4 3.1 9.5 0.5 2.6 10.1 0.4 2.4 10.6 0.4 1.2 14.0 0.3 2.8
Nose Hoover 10.0 0.3 0.3 9.9 0.5 0.1 9.6 0.6 0.1 11.5 0.3 0.5 12.5 0.7 0.5 13.0 0.5 0.5 10.6 0.9 0.8 10.6 0.9 0.7 10.1 0.7 0.7 10.2 0.5 0.4 10.2 0.7 0.5 12.9 0.4 0.4
Pehlivan Wei 4.0 0.2 0.3 8.5 0.4 0.4 10.1 0.4 0.3 11.3 0.3 0.3 10.8 0.4 0.3 12.6 0.5 0.7 9.2 0.7 1.0 10.1 0.8 1.0 9.2 0.7 0.6 10.6 0.8 1.3 9.3 0.7 1.0 12.2 0.6 1.9
Sakarya 1.5 0.2 0.3 5.6 0.5 0.2 5.0 0.3 0.2 7.5 0.2 0.3 11.1 0.4 0.2 12.5 0.4 0.4 10.2 0.9 0.4 10.0 0.6 1.4 7.7 0.9 0.9 9.1 0.4 0.8 13.1 0.8 0.8 12.6 0.4 0.7
San Um Srisuchinwong 4.1 0.2 0.2 12.9 0.4 1.0 12.7 0.5 1.6 11.4 0.5 1.0 13.3 0.5 1.2 13.0 0.6 1.1 12.7 0.7 0.9 10.4 0.7 0.9 13.4 0.6 1.2 11.7 0.7 1.5 10.7 0.7 1.6 13.1 0.4 2.4
Sprott A 2.7 0.1 0.3 8.0 0.2 0.3 8.9 0.3 0.3 6.6 0.2 0.2 11.8 0.3 0.3 12.3 0.3 1.1 10.5 0.8 1.0 8.1 0.5 1.6 10.0 0.7 0.7 10.4 0.7 1.0 9.9 0.7 0.8 12.0 0.3 3.0
Sprott B 2.4 0.2 0.3 4.3 0.2 0.8 5.6 0.4 1.6 11.7 0.5 0.2 12.7 0.4 0.6 13.2 0.7 1.7 11.8 0.8 0.8 11.5 0.6 0.6 12.3 0.8 1.2 10.6 0.7 1.4 12.5 0.8 1.8 12.6 0.4 1.5
Sprott C 2.3 0.1 0.6 11.6 0.4 1.2 12.4 0.4 0.7 11.9 0.4 1.2 13.3 0.3 0.3 13.1 0.4 2.4 10.0 0.5 4.7 9.2 0.5 3.9 10.7 0.5 5.3 9.3 0.5 4.3 12.8 0.6 2.9 13.3 0.2 2.4
Sprott D 4.0 0.2 0.3 9.8 0.6 0.1 10.3 0.4 0.1 12.9 0.3 0.2 13.4 0.5 0.4 13.4 0.4 0.1 12.2 0.7 0.2 10.9 0.6 0.2 9.9 0.5 0.6 11.5 0.6 0.6 12.9 0.7 0.2 14.0 0.2 1.4
Sprott F 3.0 0.2 0.3 9.6 0.5 0.3 9.2 0.5 0.6 12.3 0.3 0.4 12.6 0.4 0.7 12.7 0.4 0.6 9.9 0.7 1.5 10.4 0.7 1.3 10.4 0.6 1.0 12.1 0.5 1.1 11.4 0.8 1.1 13.3 0.3 0.9
Sprott G 2.3 0.1 0.7 7.0 0.6 0.1 7.7 0.5 0.1 10.0 0.3 0.3 7.2 0.1 0.4 13.0 0.3 0.6 10.9 0.9 1.2 11.4 0.6 1.3 9.2 0.8 1.4 10.8 0.8 1.7 14.1 0.8 0.7 13.1 0.2 1.3
Sprott I 2.7 0.1 0.5 9.7 0.5 0.4 11.3 0.2 0.4 9.0 0.1 0.4 11.9 0.3 0.5 13.7 0.2 1.1 10.2 0.9 2.1 9.8 0.9 3.4 8.4 0.6 2.9 9.1 0.6 2.8 12.2 0.7 1.9 14.4 0.2 6.2
Sprott L 3.4 0.1 0.4 8.0 0.5 0.3 6.8 0.1 0.3 8.2 0.2 0.1 11.6 0.3 0.2 13.0 0.2 0.1 10.5 0.9 1.2 10.4 0.8 1.8 11.4 0.9 1.2 12.1 0.7 1.0 10.8 0.8 1.3 14.2 0.3 3.1
Sprott M 3.8 0.1 0.4 9.8 0.4 0.0 10.5 0.4 0.1 12.4 0.1 0.1 12.9 0.2 0.2 14.0 0.6 0.6 9.8 0.8 1.4 10.3 0.6 1.6 9.9 0.8 1.5 11.8 0.6 1.6 11.6 0.7 1.5 14.3 0.2 3.2
Sprott O 1.3 0.1 0.9 3.6 0.1 0.3 4.6 0.0 0.5 9.7 0.1 0.4 12.4 0.4 0.4 12.9 0.4 0.9 10.8 1.0 1.3 12.6 0.9 1.3 10.4 0.8 1.6 11.5 0.9 1.1 10.7 0.8 0.6 14.0 0.4 2.0
Sprott P 5.5 0.2 0.6 10.8 0.3 2.2 11.9 0.3 1.3 12.7 0.3 1.1 12.7 0.3 1.2 13.3 0.3 2.0 7.3 0.5 3.1 10.6 0.6 4.5 8.3 0.6 4.2 10.0 0.6 5.0 11.1 0.7 2.7 12.2 0.3 5.7
Sprott Q 4.5 0.1 0.5 10.1 0.2 0.2 7.6 0.1 0.2 7.3 0.1 0.3 11.6 0.2 0.2 13.2 0.3 0.8 11.8 0.8 2.5 11.6 0.7 1.1 11.9 0.6 0.5 11.2 0.6 2.8 10.9 0.7 0.7 13.1 0.4 4.3
Sprott R 3.5 0.3 0.4 9.3 0.8 0.1 6.7 0.5 0.2 11.0 0.5 0.4 12.6 0.5 0.6 13.1 0.4 0.3 10.6 0.8 0.4 10.4 0.5 0.3 9.6 0.7 0.8 10.5 0.7 0.8 11.8 0.8 0.8 13.8 0.3 1.3
Sprott S 5.3 0.1 0.6 10.1 0.9 1.7 9.9 0.6 0.7 12.3 0.2 0.9 12.6 0.4 0.4 13.1 0.3 1.1 10.6 0.7 0.6 9.7 0.8 3.5 11.0 0.6 1.8 10.4 0.4 3.7 11.8 0.8 1.8 11.6 0.3 4.6
Sprott Torus 3.2 0.2 0.1 5.4 0.7 0.4 10.4 0.7 0.1 14.3 0.6 0.2 14.5 0.6 0.4 13.8 0.7 0.3 6.2 0.8 2.9 8.3 0.6 5.2 5.3 0.8 5.4 5.8 0.7 4.1 13.6 0.8 1.0 15.5 0.3 2.1
Thomas 1.4 0.3 2.1 10.6 0.5 3.8 10.1 0.8 3.9 9.9 0.3 3.7 13.9 0.5 2.6 13.9 0.7 2.8 12.6 0.8 1.9 12.6 0.8 2.0 12.8 0.8 2.0 12.9 0.8 2.0 10.7 0.5 3.0 13.3 0.2 2.6
Thomas Labyrinth 4.4 0.1 1.8 12.5 0.2 2.8 12.5 0.3 2.9 12.4 0.4 3.1 12.5 0.4 2.9 12.5 0.3 3.8 11.4 0.5 2.7 12.5 0.5 2.2 11.9 0.5 2.5 11.7 0.5 2.1 12.3 0.5 3.1 12.4 0.2 1.8
Torus 11.5 0.9 2.7 9.8 0.3 0.1 9.4 0.2 0.2 10.9 0.3 0.2 12.3 0.3 0.3 12.4 0.5 0.3 12.1 0.9 1.9 10.5 1.0 2.0 11.6 0.9 1.8 11.9 0.5 1.9 9.1 0.8 1.9 12.6 0.5 3.2
Tsucs 2 5.7 0.6 0.4 11.6 0.3 0.1 10.3 0.5 0.2 10.6 0.3 0.3 11.6 0.6 0.2 13.8 0.7 0.9 10.6 0.7 0.4 11.2 0.8 0.4 8.2 0.7 0.4 10.5 0.8 0.4 11.1 0.7 0.4 8.4 0.4 0.5
Vallise El Nino 1.2 0.1 0.3 8.4 0.3 0.6 11.2 0.5 1.2 12.5 0.3 1.3 12.7 0.5 1.9 13.5 0.6 3.2 12.3 0.7 2.5 13.4 0.5 2.7 12.3 0.7 3.3 12.6 0.5 1.2 13.9 0.8 2.9 13.7 0.3 3.7
Wang Sun 2.8 0.2 0.1 13.2 0.5 0.5 12.2 0.5 0.2 11.8 0.5 0.5 11.1 0.4 0.5 14.6 0.7 1.2 13.9 0.7 0.9 10.9 0.7 0.9 10.8 0.6 1.1 10.9 0.7 0.8 13.3 0.7 0.6 11.1 0.4 0.8
Winsmi Reduced 6.5 0.4 0.2 11.4 0.4 0.1 10.7 0.5 0.1 13.5 0.6 1.1 13.7 0.6 0.1 14.0 0.4 1.3 11.9 0.9 3.7 12.2 0.8 4.0 10.5 0.6 3.6 9.4 0.3 2.3 12.0 0.8 3.2 14.2 0.2 3.3
Yu Wang 5.5 0.2 0.6 9.8 0.7 0.8 8.4 0.3 0.3 12.4 0.3 0.2 12.4 0.5 0.7 13.0 0.6 0.9 12.4 0.9 3.6 12.5 0.7 3.2 12.8 0.8 3.3 11.9 0.9 3.9 12.5 0.8 2.3 13.5 0.4 2.9
Yu Wang 2 3.8 0.0 0.1 9.6 0.5 0.1 10.0 0.6 0.1 12.4 0.3 0.2 12.2 0.3 0.1 14.1 0.7 0.2 9.9 0.2 0.2 9.5 0.7 0.4 9.9 0.6 0.6 10.4 0.4 0.3 10.6 0.7 0.6 12.6 0.2 0.4
Zhou Chen 4.9 0.3 0.2 11.1 0.3 0.5 5.9 0.4 0.2 11.6 0.3 0.9 14.2 0.4 0.7 10.3 0.3 0.8 12.8 0.4 0.7 12.7 0.5 0.9 13.5 0.4 1.0 12.9 0.5 1.2 13.1 0.5 0.6 13.2 0.2 2.1
Table 3: Performance comparison for zero-shot inference across all DS in the dataset from [29] using $T_{C}=512$. Best value in red, second best in blue.
System DynaMix Chronos-t5-large Chronos-t5-base Chronos-t5-small Chronos-t5-mini Chronos-t5-tiny Chronos-bolt-base Chronos-bolt-small Chronos-bolt-mini Chronos-bolt-tiny Mamba4Cast TTM
$D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE $D_{\text{stsp}}$ $D_{H}$ MAE
Atmospheric Regime 8.4 0.6 2.8 9.7 0.5 0.4 11.1 0.4 0.6 10.5 0.6 0.6 10.1 0.5 1.3 12.0 0.4 0.5 11.9 0.9 3.6 12.0 0.9 3.8 13.5 0.9 2.7 12.7 0.8 4.9 11.3 0.9 1.2 14.2 0.6 3.6
BeerRNN 11.7 0.8 4.7 8.0 0.5 0.5 3.3 0.2 0.4 9.5 0.2 0.2 7.3 0.4 0.2 11.9 0.4 0.3 12.3 0.9 1.4 12.5 0.9 2.2 12.1 0.9 1.8 12.6 0.8 1.4 13.0 0.9 1.9 14.1 0.6 4.1
Bouali 4.1 0.1 0.2 8.3 0.5 0.0 7.3 0.3 0.1 8.6 0.1 0.0 11.1 0.3 0.1 13.2 0.5 0.1 10.6 0.8 0.9 9.6 0.7 1.1 12.2 0.7 1.2 12.3 0.8 0.8 12.9 0.8 1.5 13.4 0.4 1.1
Burke Shaw 4.1 0.2 0.6 10.8 0.5 1.5 9.0 0.5 2.4 13.0 0.4 2.0 13.2 0.4 1.9 13.3 0.6 1.8 11.5 0.8 0.6 12.1 0.6 1.0 12.2 0.7 1.2 10.5 0.7 2.9 12.2 0.9 1.3 13.7 0.5 2.6
Ca Two Plus Quasip. 2.3 0.1 0.1 7.2 0.1 0.1 8.3 0.1 0.1 9.2 0.4 0.1 13.1 0.4 0.1 13.2 0.4 0.2 11.3 0.9 0.8 11.6 0.8 1.2 12.8 0.6 0.3 8.9 0.6 1.6 11.1 0.8 1.6 13.8 0.7 1.6
Chen Lee 3.9 0.4 0.3 4.9 0.3 0.9 4.0 0.5 1.6 4.8 0.3 0.4 4.6 0.3 1.0 8.9 0.7 1.3 7.2 0.9 3.6 7.9 0.7 2.1 8.0 0.7 1.7 4.8 0.5 1.4 11.0 0.9 1.2 6.4 0.5 4.0
Chua 5.9 0.1 0.1 7.4 0.2 0.2 5.9 0.2 0.2 7.6 0.1 0.2 11.7 0.2 0.1 13.1 0.3 0.4 11.8 0.6 1.0 12.2 0.5 1.3 11.7 0.6 0.6 12.0 0.6 1.4 11.3 0.5 1.5 13.4 0.4 0.8
Coullet 1.2 0.0 0.4 4.9 0.0 0.2 0.7 0.0 0.1 2.8 0.0 0.1 7.5 0.2 0.0 14.6 0.4 0.3 13.0 0.8 1.7 11.2 0.8 2.4 11.8 0.8 2.3 11.6 0.7 2.4 12.3 0.8 1.7 14.9 0.5 1.9
Dadras 4.1 0.2 1.5 7.7 0.6 0.9 4.7 0.5 1.2 13.1 0.5 0.5 12.1 0.4 0.3 10.2 0.7 1.0 12.7 0.8 2.6 12.9 0.7 2.4 13.0 0.8 2.7 8.8 0.8 4.2 11.9 0.8 2.5 12.8 0.4 2.8
Dequan Li 12.7 0.5 0.3 11.1 0.3 0.2 11.0 0.4 0.4 12.1 0.3 0.3 12.1 0.2 0.5 13.0 0.3 1.5 11.7 0.4 1.1 11.3 0.4 1.0 9.6 0.4 0.7 11.0 0.4 0.4 10.7 0.4 0.6 13.5 0.4 2.5
Finance 3.2 0.1 0.4 8.5 0.6 0.9 8.6 0.4 0.7 9.1 0.4 0.3 12.5 0.3 0.1 13.8 0.4 1.4 11.6 0.9 1.8 10.7 0.8 2.0 13.4 0.8 1.5 11.8 0.8 1.1 11.1 0.9 0.8 13.7 0.5 1.3
Genesio Tesi 8.2 0.2 0.3 7.9 0.4 0.7 8.5 0.2 0.7 7.7 0.2 0.7 9.4 0.2 0.8 13.0 0.3 1.6 10.4 0.7 0.7 10.6 0.7 0.8 10.3 0.8 1.0 9.6 0.8 1.5 11.0 0.9 0.9 12.6 0.6 4.1
Guckenheimer Holmes 9.6 0.1 0.2 11.8 0.1 0.4 12.0 0.1 0.4 12.6 0.3 0.2 12.5 0.4 0.6 12.1 0.4 0.2 9.6 0.7 0.2 9.6 0.7 0.4 9.4 0.6 0.2 11.8 0.6 0.4 10.2 0.7 0.5 12.7 0.5 0.6
Hadley 11.3 0.4 0.6 10.7 0.5 2.3 9.7 0.5 1.3 12.5 0.3 1.1 12.6 0.6 2.1 12.6 0.4 3.2 11.4 0.8 3.2 11.9 0.7 3.3 11.9 0.6 2.9 11.5 0.8 3.4 11.4 0.8 3.5 12.5 0.5 3.6
Halvorsen 4.3 0.3 1.0 9.7 0.6 0.3 8.2 0.2 0.8 7.5 0.2 0.8 13.1 0.2 0.6 13.3 0.4 1.3 10.0 0.8 4.7 10.3 0.7 4.4 9.7 0.8 4.4 10.9 0.7 4.3 11.3 0.9 3.4 13.4 0.5 3.5
Hindmarsh Rose 6.5 0.4 0.5 9.8 0.2 0.1 4.9 0.1 0.0 9.6 0.2 0.0 12.7 0.6 0.2 13.7 0.6 0.5 12.9 0.7 1.1 14.3 0.8 1.3 12.9 0.8 1.1 14.0 0.8 1.4 13.2 0.8 0.6 14.6 0.4 1.1
Isothermal Chemical 2.0 0.2 0.3 2.5 0.1 0.2 5.4 0.1 0.1 8.3 0.2 0.1 7.0 0.4 0.1 11.6 0.5 0.2 11.7 0.9 0.6 11.7 0.9 1.0 10.8 0.8 0.8 12.5 0.9 0.8 10.1 0.9 0.9 15.6 0.6 2.5
Itik Banks tumor 5.7 0.2 0.1 8.2 0.5 0.1 10.8 0.6 0.2 9.8 0.5 0.1 11.8 0.6 0.2 13.3 0.6 0.1 10.8 0.6 0.9 11.2 0.8 1.7 11.9 0.9 0.9 12.0 0.8 2.4 12.0 0.9 1.4 12.9 0.7 1.5
Jerk Circuit 4.3 0.3 0.4 8.8 0.6 1.1 9.1 0.4 0.8 10.6 0.3 1.0 11.7 0.3 0.2 9.8 0.6 1.1 12.5 0.9 1.9 13.5 0.9 1.6 12.9 0.9 0.7 13.1 1.0 1.5 10.3 0.8 0.5 14.1 0.5 3.1
Laser 4.5 0.2 0.5 7.5 0.5 5.0 8.2 0.3 1.9 11.5 0.2 0.6 12.8 0.2 0.7 13.9 0.6 1.3 12.6 0.7 2.4 12.8 0.5 1.7 13.0 0.7 1.3 12.8 0.7 2.7 11.8 0.8 2.4 13.3 0.5 2.2
Lorenz 3.4 0.2 0.6 9.8 0.6 0.5 8.5 0.3 0.6 13.0 0.4 0.7 12.6 0.7 0.8 13.3 0.6 0.3 12.8 0.8 1.0 13.0 0.7 1.9 13.2 0.6 1.4 13.0 0.6 1.3 11.8 0.8 1.3 13.4 0.5 1.1
Lorenz 84 7.9 0.2 1.6 10.7 0.5 3.3 11.0 0.4 4.4 12.6 0.4 4.8 11.0 0.4 0.8 13.2 0.6 1.8 11.8 0.9 2.0 12.2 0.9 2.6 11.5 0.9 2.4 12.3 0.9 2.5 11.2 0.9 2.2 13.4 0.6 2.9
Lu Chen 9.5 0.8 1.4 5.3 0.2 1.0 9.0 0.3 1.3 8.9 0.2 0.2 9.7 0.3 0.8 8.3 0.3 1.0 12.2 1.0 1.0 13.8 0.9 1.4 12.2 0.9 1.7 13.1 1.0 0.9 10.1 0.8 1.4 14.2 0.5 2.2
Moore Spiegel 9.2 0.6 0.9 9.1 0.6 0.9 10.9 0.7 0.9 12.5 0.5 0.8 11.5 0.6 0.8 11.3 0.4 0.9 11.4 0.9 2.6 11.7 0.8 3.6 12.7 0.9 3.6 12.0 0.8 3.3 10.5 0.9 4.6 13.1 0.7 4.0
Multi Chua 12.5 0.5 0.2 12.0 0.6 0.5 11.9 0.6 0.3 13.4 0.4 0.5 13.0 0.4 0.3 13.9 0.4 0.4 12.5 0.5 0.5 12.6 0.5 0.4 13.6 0.5 0.3 13.0 0.5 0.5 10.7 0.5 0.1 13.5 0.5 0.8
Nose Hoover 11.4 0.5 0.5 10.2 0.6 0.4 10.8 0.7 0.6 10.8 0.6 0.3 12.9 0.8 0.6 13.3 0.5 0.6 11.2 0.9 1.6 11.3 0.9 1.4 12.7 0.7 2.2 11.0 0.9 2.0 10.5 0.9 2.3 13.3 0.5 3.1
Pehlivan Wei 4.8 0.2 3.1 4.6 0.1 0.4 6.8 0.4 0.5 9.0 0.3 0.6 10.7 0.3 0.9 12.4 0.5 1.5 11.1 0.9 5.8 12.0 0.6 6.1 12.6 0.8 6.1 12.3 0.8 5.5 10.0 0.9 2.3 12.7 0.4 6.3
Sakarya 3.1 0.3 1.0 2.6 0.2 0.5 5.0 0.3 0.5 8.2 0.3 0.5 10.2 0.7 1.0 12.9 0.8 1.8 10.7 0.9 1.5 12.7 0.9 1.9 11.8 0.8 2.0 12.8 0.8 2.1 12.4 0.7 1.4 14.0 0.5 2.9
San Um Srisuchinwong 5.5 0.1 0.5 5.5 0.2 1.5 10.5 0.8 1.5 12.6 0.4 1.5 13.0 0.5 1.4 13.1 0.6 0.7 11.8 0.8 1.4 13.0 0.7 0.3 12.0 0.6 1.1 11.9 0.7 0.6 10.7 0.5 1.7 12.6 0.5 0.5
Sprott A 4.9 0.1 0.6 9.3 0.5 0.2 8.5 0.4 0.3 7.5 0.3 0.3 11.7 0.3 0.6 12.6 0.5 0.8 10.7 0.9 2.9 11.8 0.7 2.9 11.5 0.8 3.1 11.1 0.7 3.5 9.8 0.9 0.6 12.4 0.6 3.5
Sprott B 2.6 0.2 2.1 9.2 0.5 1.1 11.0 0.6 1.3 13.2 0.5 2.0 13.4 0.6 0.9 13.4 0.7 2.3 12.6 0.8 1.2 13.2 0.8 1.9 12.5 0.8 2.0 12.6 0.7 1.8 13.9 0.8 1.9 12.9 0.6 1.5
Sprott C 2.9 0.2 0.3 8.5 0.6 1.2 12.0 0.3 1.2 13.3 0.4 0.5 12.6 0.4 0.6 11.8 0.5 2.4 12.4 0.7 2.2 13.0 0.6 2.1 11.9 0.6 2.9 11.8 0.6 2.1 13.3 0.7 2.2 13.5 0.5 2.9
Sprott D 3.8 0.1 0.3 7.6 0.2 0.1 5.5 0.1 0.1 5.5 0.2 0.1 12.9 0.4 0.5 13.1 0.5 0.6 12.1 0.7 0.5 12.0 0.8 3.1 12.1 0.7 1.8 12.0 0.6 2.0 12.9 0.7 0.7 13.8 0.5 1.1
Sprott F 4.9 0.3 1.1 10.8 0.4 0.9 8.9 0.5 0.7 12.6 0.3 0.5 13.0 0.5 0.5 13.1 0.4 0.4 12.3 0.9 1.8 12.2 0.6 2.1 13.2 0.8 2.0 12.4 0.7 2.4 11.8 0.9 1.0 13.4 0.6 1.3
Sprott G 2.4 0.0 0.2 4.2 0.1 0.1 6.2 0.2 0.1 7.1 0.2 0.2 7.7 0.3 0.7 12.9 0.3 0.6 12.3 0.8 3.5 12.4 0.7 3.9 12.9 0.8 2.9 11.2 0.9 4.4 11.0 0.9 2.3 14.0 0.5 1.6
Sprott I 3.0 0.2 0.6 8.6 0.5 0.1 10.8 0.1 0.1 7.4 0.3 0.3 10.1 0.3 0.3 13.1 0.4 0.2 12.1 0.9 1.1 11.9 0.8 2.9 9.4 0.9 1.0 13.4 0.9 2.2 10.9 0.9 1.0 14.4 0.6 2.7
Sprott L 2.7 0.2 1.4 9.2 0.3 0.2 8.2 0.1 0.3 10.2 0.2 0.2 11.3 0.2 0.3 13.6 0.4 0.2 12.3 0.9 1.0 12.3 0.9 0.4 13.3 0.9 1.6 13.1 0.9 2.0 12.2 0.8 0.5 14.2 0.5 3.1
Sprott M 3.4 0.2 0.4 10.9 0.5 0.1 12.4 0.4 0.1 12.8 0.3 0.3 13.5 0.3 0.3 14.0 0.8 0.5 12.1 0.8 0.9 11.1 0.7 1.3 13.0 0.8 0.9 12.3 0.8 0.8 10.7 0.9 0.8 14.3 0.6 2.3
Sprott O 3.5 0.2 0.4 9.2 0.2 0.4 6.7 0.1 0.4 8.1 0.2 0.6 11.6 0.4 0.4 13.2 0.3 0.5 11.8 1.0 0.9 12.7 0.8 1.6 11.4 1.0 1.8 12.6 1.0 3.9 10.4 0.9 0.9 14.0 0.7 3.5
Sprott P 5.7 0.2 1.2 9.2 0.6 0.8 10.5 0.4 0.5 11.1 0.8 1.0 13.3 0.4 0.7 13.5 0.6 1.6 9.6 0.8 4.0 7.6 0.6 3.3 8.7 0.8 3.4 9.6 0.8 3.0 10.8 0.8 1.5 10.4 0.5 1.8
Sprott Q 5.8 0.2 1.1 9.5 0.5 0.5 9.4 0.5 0.4 10.0 0.7 0.5 12.9 0.5 1.3 12.6 0.3 1.3 11.1 0.9 0.8 13.0 0.7 1.8 12.5 0.7 0.6 11.5 0.8 1.7 9.7 0.9 0.4 13.2 0.4 4.0
Sprott R 5.1 0.3 0.5 6.9 0.4 0.1 6.7 0.3 0.1 6.5 0.2 0.0 12.5 0.3 0.3 14.2 0.5 0.4 11.0 0.9 1.4 8.8 0.7 1.0 10.6 0.9 1.0 9.4 0.9 0.9 11.4 0.9 1.7 13.8 0.6 2.3
Sprott S 2.9 0.0 0.3 9.6 0.4 0.3 10.7 0.2 0.2 11.7 0.3 0.3 13.0 0.4 0.3 13.3 0.4 0.7 10.2 0.9 0.6 10.2 0.8 1.6 11.2 0.9 1.7 11.0 0.8 2.0 10.7 0.9 1.9 12.8 0.5 4.0
Sprott Torus 3.6 0.2 0.2 6.8 0.6 0.2 12.0 0.5 0.2 13.0 0.5 0.3 14.3 0.7 0.3 14.7 0.6 0.7 7.0 0.9 0.6 12.1 0.4 0.8 7.9 0.9 0.6 12.5 0.7 0.1 12.2 0.9 0.3 14.9 0.5 1.0
Thomas 2.3 0.2 6.7 12.4 0.6 4.1 12.1 0.3 3.6 13.4 0.2 2.5 13.4 0.3 1.9 13.8 0.2 1.1 12.3 0.8 4.9 12.0 0.8 6.6 13.7 0.7 4.9 13.1 0.7 5.5 11.9 0.8 4.9 13.6 0.4 5.4
Thomas Labyrinth 6.5 0.1 5.2 12.5 0.2 4.8 9.8 0.4 5.6 11.8 0.3 3.4 12.2 0.3 4.6 12.5 0.4 4.0 11.5 0.6 4.4 12.3 0.6 4.2 11.5 0.6 4.2 11.5 0.6 3.8 12.5 0.6 3.3 12.5 0.3 2.6
Torus 10.9 0.9 2.3 12.4 0.4 0.1 11.9 0.2 0.1 12.0 0.2 0.2 12.9 0.5 1.0 12.9 0.7 1.5 12.0 1.0 2.1 12.6 0.8 2.3 12.6 1.0 2.0 12.4 1.0 3.1 8.7 0.9 1.7 13.1 0.6 4.6
Tsucs 2 7.2 0.6 1.7 10.9 0.7 1.0 10.3 0.8 1.2 10.0 0.2 1.6 12.0 0.6 1.7 13.3 0.7 2.2 8.5 0.8 2.2 12.3 0.8 2.2 7.5 0.7 2.4 9.4 0.7 2.3 11.5 0.8 2.6 9.8 0.5 2.1
Vallise El Nino 1.8 0.2 0.4 10.1 0.4 0.3 5.5 0.2 0.2 7.4 0.3 0.2 12.7 0.5 0.5 13.4 0.6 1.7 11.6 0.8 1.8 12.5 0.8 1.0 13.6 0.7 1.1 12.4 0.8 1.9 13.6 0.7 1.3 13.7 0.6 2.1
Wang Sun 5.3 0.2 0.7 11.8 0.5 1.2 10.3 0.4 0.4 13.2 0.5 2.0 10.4 0.5 0.4 10.5 0.7 0.7 9.3 0.7 1.9 11.9 0.7 1.8 11.2 0.7 1.8 9.9 0.6 1.7 13.3 0.7 1.5 12.6 0.5 1.0
Winsmi Reduced 5.0 0.3 1.4 8.9 0.4 1.8 11.2 0.4 1.6 13.6 0.4 1.6 13.6 0.5 1.5 14.0 0.7 1.0 11.7 0.9 2.1 11.7 0.8 2.4 11.9 0.9 1.4 10.4 0.8 2.7 11.4 0.9 0.9 14.3 0.6 2.1
Yu Wang 6.4 0.3 3.2 9.8 0.7 3.3 11.3 0.6 3.4 10.3 0.6 3.5 8.1 0.6 3.3 12.8 0.6 3.2 12.3 0.9 7.6 12.0 0.8 5.7 12.8 0.7 6.6 12.8 0.8 7.2 12.0 0.7 5.0 13.7 0.5 4.6
Yu Wang 2 3.4 0.1 0.3 10.0 0.2 0.1 10.6 0.4 0.2 12.7 0.2 0.1 12.2 0.2 0.5 13.8 0.4 0.5 10.8 0.9 1.4 9.4 0.8 2.5 8.6 0.7 1.6 10.0 0.7 1.4 11.6 0.9 1.8 10.7 0.3 1.2
Zhou Chen 10.0 0.4 0.2 10.4 0.4 1.3 13.7 0.5 3.0 12.7 0.4 0.9 12.8 0.3 0.4 12.8 0.4 0.9 13.2 0.5 1.6 11.6 0.5 2.0 13.5 0.5 1.8 9.7 0.5 2.2 12.8 0.6 2.7 11.6 0.4 2.9